OpenStudy (anonymous):

Why is is that as limit of x goes to infinity, that 1/arctan(x) becomes 1/(pi/2).....

OpenStudy (anonymous):

because \[lim_{x\to\infty}arctan(x)=\frac{\pi}{2} \]

OpenStudy (anonymous):

and that is because as \[lim_{x\to\frac{\pi}{2}} tan(x)=\infty\]

OpenStudy (anonymous):

thanks, it all makes sense now